Yield to Maturity and Current Yield

You just purchased a bond that matures in 4 years. The bond has a face value of $1,000 and has an 9% annual coupon. The bond has a current yield of 7.63%. What is the bond's yield to maturity? Round your answer to two decimal places.

_____________%
Answer #1

Coupon =9%*1000 =90
Current Yield =Coupon/Current Price
Price =90/7.63% =1179.5544
Par Value =1000
YTM using excel formula =RATE(4,90,-1179.5544,1000) =4.05%